# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 3.4. Fixed-parameter ABMN asymptotics
(1,2). From $c(x)=1/\gamma(\kappa,\rho,\beta)$ and (9), we find that
@@
We have that $c_{-j}(x)=c\big(s_{-j}(x)\big)=1/\gamma(\kappa,\rho,\beta_{j})$, with $\beta_{-j}$-asymptotics offered by Lemma 3.5. When $\rho=1$,
@@
$\frac{m_{-(i+1),-i}}{m_{-1,0}}$ $=$ $\prod_{j=1}^{i}\left(c_{j}-1\right)^{-1}=\prod_{j=1}^{i}\left(\frac{1-\kappa}{1+\kappa}\right)\left(1+\frac{1-\rho}{2\rho^{2}}j^{-1}\Big(1+O(j^{-1})\Big)\right)$
@@
Noting $\big(M_{-i}/N_{-i}\big)^{\rho}=\beta_{-i}^{\rho}\leq C/i$ offers a simplified asymptotic formula for $a_{-i}$, namely
@@
for a suitable constant $\chi=\chi(\rho,\sigma)$. From $M_{i}=m_{-(i+1),-i}+m_{-i,-(i-1)}$ and $m_{-(i+1),-i}\ll m_{-i,-(i-1)}$, we see that
@@
4. Nash equilibria and the ABMN equations
Here we prove Theorem 1.8 on Nash-ABMN equivalence. The forward implication $(1)\!\!\!\!\implies\!\!\!\!(2)$ is proved in the first four subsections, the reverse in the fifth. The derivations follow the template given by the proof of the counterpart [26, Theorem $2.6$] in [26, Chapter $4$], with some substantial changes.
In the forward-implication proof, some arguments are new and others closely follow counterparts in [26, Chapter $4$]. To make our presentation self-contained while indicating where the overlap lies, the first three subsections use the convention that Proof denotes the start of an argument with substantial new elements, while Derivation indicates one that is close to one in [26]. No lack of rigour should be inferred from use of the latter label, though we have sometimes opted for a more verbal style of presentation of such arguments. A different approach has been adopted for the reverse implication, as we explain in Section 4.5.
@@ 4.1. Escape is almost certain at a time-invariant Nash equilibrium
Recall the payoff notation (4). A strategy pair $(S_{-},S_{+})\in\mathcal{S}^{2}$ is said to have finite mean costs if neither $\mathbb{E}^{k}_{S_{-},S_{+}}[P_{-}]$ nor $\mathbb{E}^{k}_{S_{-},S_{+}}[P_{+}]$ equals minus infinity, for any $k\in\mathbb{Z}$.
Let $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}$. Denote $b_{i}=S_{-}(i,j)$ and $a_{i}=S_{+}(i,j)$ for $(i,j)\in\mathbb{Z}\times\mathbb{N}_{+}$ here also (without supposing $(S_{-},S_{+})\in\mathcal{N}_{\kappa,\rho}$). The idle zone $\mathcal{I}$ is set equal to $\big\{j\in\mathbb{Z}:a_{j}=b_{j}=0\big\}$.
@@
Derivation. At each turn whose index lies in $\llbracket i,j-1\rrbracket$, the counter lies in the idle zone and no stakes are offered. The counter thus evolves as a symmetric simple random walk: on flip moves, by definition; on stake moves, by the zero-stake rule given in Section 1.3. ∎
@@
$\textrm{Rocket}_{\psi}^{\leftarrow i}(j)\,=\,\psi^{i-j+1}{\bf 1}_{j\leq i}\,\,\,,\,\,\,j\in\mathbb{Z}\,$,
@@
    $\Big(\frac{1-\kappa}{1+\kappa}\Big)^{k+1}\big(m_{\infty}-m_{*}\big)+\bigg(\psi^{k}+\Big(\frac{1-\kappa}{1+\kappa}\Big)^{k+1}\bigg)\kappa^{-1}\bigg(\frac{\psi}{1-\psi}+\frac{1-\kappa}{(1+\kappa)\big(1-\frac{1-\kappa}{(1+\kappa)\psi}\big)}\bigg)$ (27)
@@
    If $\textrm{Left}(S_{-})$ and $\textrm{Left}(S_{+})$ exceed -$\infty$ and $i\in\mathbb{Z}$ is less than their minimum, then, provided that the quantity given by replacing $m_{\infty}-m_{*}$ by $n_{-\infty}-n_{*}$ in (27) is strictly less than $n_{\infty,-\infty}$, we have that $\mathbb{E}^{i-k}_{\textrm{Rocket}_{\psi}^{\leftarrow i}(S_{-}),S_{+}}[P_{-}]>\mathbb{E}^{i-k}_{S_{-},S_{+}}[P_{-}]$.
@@
$\mathbb{E}^{i+k}_{S_{-},\textrm{Rocket}_{\psi}^{i\rightarrow}}\big[C_{+}[0,\tau_{i-1})\big]$ $=$ $\sum_{j=i}^{\infty}\mathbb{E}_{i+k}\big[\#^{[0,\tau_{i-1})}_{j}(Z)\big]\psi^{j-i+1}$ (28)
@@
$\mathbb{E}^{i+k}_{S_{-},\textrm{Rocket}_{\psi}^{i\rightarrow}}\big[C_{+}[\tau_{i-1},\infty)\big]\,\leq\,\big(\tfrac{1-\kappa}{1+\kappa}\big)^{k+1}\kappa^{-1}\bigg(\frac{\psi}{1-\psi}+\frac{1-\kappa}{(1+\kappa)\big(1-\frac{1-\kappa}{(1+\kappa)\psi}\big)}\bigg)\,.$ (29)
From (28) and (29), we find that
$\mathbb{E}^{i+k}_{S_{-},\textrm{Rocket}_{\psi}^{i\rightarrow}}\big[C_{+}\big]\,\leq\,\Big(\psi^{k}+\big(\tfrac{1-\kappa}{1+\kappa}\big)^{k+1}\Big)\kappa^{-1}\bigg(\frac{\psi}{1-\psi}+\frac{1-\kappa}{(1+\kappa)\big(1-\frac{1-\kappa}{(1+\kappa)\psi}\big)}\bigg)\,.$ (30)
@@
while the second is at most the right-hand side of (30). Hence, the hypothesis on $k$ expressed in terms of (27) implies that $\mathbb{E}^{i+k}_{S_{-},\textrm{Rocket}_{\psi}^{i\rightarrow}}\big[P_{+}\big]-\mathbb{E}^{i+k}_{S_{-},S_{+}}\big[P_{+}\big]$ is strictly positive, as we seek to show in proving Lemma 4.4(1).
@@
Let $(S_{-}^{\prime},S_{+})$ be a left strengthening of $(S_{-},S_{+})$ For $i\in\mathbb{Z}$, there is a coupling of gameplays $X,X^{\prime}:\mathbb{N}\to\mathbb{Z}$ under $\mathbb{P}^{i}_{S_{-},S_{+}}$ such that $X^{\prime}(j)\leq X(j)$ for $j\in\mathbb{N}$ almost surely. Couplings with the evidently needed direction for the bounds exist for each of the three other variations.
@@
    If $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}$ satisfies $\textrm{Left}(S_{-})>$-$\infty$ and $\textrm{Left}(S_{+})=$-$\infty$, let $i\in\mathbb{Z}$ satisfy $S_{+}(i,1)>0$ and $S_{-}(j,1)=0$ for $j\in(-\infty,i-1\rrbracket$. Then $\mathbb{P}^{i}_{S_{-},S_{+}}(E_{-})$ equals zero.
@@ 3. $(3)$
In the remaining case, $\textrm{Left}(S_{-})>$-$\infty$ and $\textrm{Left}(S_{+})=$-$\infty$. The pair $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}\cap\mathcal{N}_{\kappa,\rho}$ is non-zero, because $S_{+}$ is; it has finite mean costs by Lemma 4.7(1). Thus $\mathbb{P}^{i}_{S_{-},S_{+}}(E^{c})=0$ by Proposition 4.3. Select $i\in\mathbb{Z}$ such that $S_{+}(i,1)>0$ and $S_{-}(j,1)=0$ for $j\in(-\infty,i\rrbracket$. Lemma 4.7(2) implies that $\mathbb{P}^{i}_{S_{-},S_{+}}(E_{-})=0$. Thus, $\mathbb{P}^{i}_{S_{-},S_{+}}(E_{+})=1$, so that $T_{+}$ equals $m_{\infty}$ almost surely. If Maxine plays a strategy $\hat{S}_{+}$ formed from $S_{+}$ by reducing the stake she offers at $i$ by a factor of one-half, then gameplay $X:\mathbb{N}\to\mathbb{Z}$ is equal in law under $\mathbb{P}^{i}_{S_{-},S_{+}}$ and $\mathbb{P}^{i}_{S_{-},\hat{S}_{+}}$; $T_{+}=m_{\infty}$ almost surely under both laws; but Maxine’s running cost is almost surely less under $\mathbb{P}^{i}_{S_{-},\hat{S}_{+}}$ than it is under $\mathbb{P}^{i}_{S_{-},S_{+}}$, because the first cost, incurred at site $i$, is lower. Thus, $\mathbb{E}^{i}_{\hat{S}_{-},S_{+}}[P_{+}]>\mathbb{E}^{i}_{S_{-},S_{+}}[P_{+}]$.
@@ 4.2. A Nash component wins against zero
The presence of flip moves, when $\kappa\in(0,1)$, makes the proposition non-trivial, as we now explain. In the setup in question, $S_{-}$ is known to be wide by Lemma 4.7(3); so Mina offers positive stakes at an infinite set $K$ of integer sites. When $\kappa$ equals one (as it is in [26]), so that every move is stake, this is enough to reach the desired conclusion that left escape $E_{-}$ is almost certain starting from given $i\in\mathbb{Z}$. Indeed, when $X$ visits $K$, a left move is assured; while at sites in $\mathbb{Z}\setminus K$, no stakes are offered by either player, and the next move has equal chance of being left or right, according to the rule for zero stakes given in Section 1.3. It is easily seen that this dynamics forces the counter leftward, through a sequence of one-way locks. However, when $\kappa\in(0,1)$, flip moves occur with probability $1-\kappa$; so, when $X$ visits $K$, the next move is left with probability $(1+\kappa)/2$. The counter thus evolves as a symmetric simple random walk on $\mathbb{Z}\setminus K$, with moves biased to the left by a uniform amount on visits to $K$. Although $K$ is infinite (since $S_{-}$ is wide), this set could in principle be arbitrarily sparse; in which case, this dynamics will not realize left escape $E_{-}$ for some (or indeed all) starting points.
@@
We will in fact prove the stronger assertion that, when $(S_{-},S_{+})\in\mathcal{N}_{\kappa,\rho}$ with $S_{-}\in\mathcal{S}_{0}$ satisfies $\mathbb{P}^{i}_{(S_{-},0)}(E_{-})=0$ for some $i\in\mathbb{Z}$ such that $S_{-}(1)>0$ when $X_{0}=i$, then $(S_{-},S_{+})\not\in\mathcal{N}_{\kappa,\rho}$. Fixing such an $i$, we will show that
$\mathbb{E}^{i}_{(0,S_{+})}[P_{-}]>\mathbb{E}^{i}_{(S_{-},S_{+})}[P_{-}]\,:$ (31)
it is in Mina’s interests to play the zero strategy, rather than $S_{-}$, against Maxine’s $S_{+}$, when play starts at $i$. Naturally, (31) implies that $(S_{-},S_{+})\not\in\mathcal{N}_{\kappa,\rho}$, so proving (31) is enough.
Preparing to show (31), note that
$\mathbb{P}^{i}_{(S_{-},S_{+})}(E_{-})=\mathbb{P}^{i}_{(0,S_{+})}(E_{-})=0\,.$ (32)
Indeed, $(S_{-},0)\to(S_{-},S_{+})$ is a right strengthening and $(S_{-},S_{+})\to(0,S_{+})$ is a left weakening, so (32) follows from $\mathbb{P}^{i}_{(S_{-},0)}(E_{-})=0$ and Lemma 4.6.
Why may we expect (31) to hold? In other words, why would Mina switch from $S_{-}$ to 0 against $S_{+}$? That $\mathbb{P}^{i}_{(S_{-},S_{+})}(E_{-})$ is zero makes Mina’s motivation simple: $S_{-}$ is not working out for her, because her victory $E_{-}$ never happens. By switching to 0, she will save on running costs. As for terminal receipts, these are split between non-escape $E^{c}$ and right escape $E_{+}$ when she plays $S_{-}$. By playing 0 instead, Mina will cease to exert any left pressure, so, in an instance of right strengthening and monotonicity. any change to this split will take the form of a rightward move of probability mass from $E^{c}$ to $E_{+}$. But that would help Mina, because $E^{c}$ is the worse outcome for her in the sense that $n_{*}\leq n_{\infty}$. To record these inferences symbolically,
@@
where the first equality is due to absence of running cost for Mina when she plays zero; the second equality crucially invokes $\mathbb{P}^{i}_{(S_{-},0)}(E_{-})=0$; the first inequality is due to the (32)-consequences
@@
and the monotonicity deduction $\mathbb{P}^{i}_{(0,S_{+})}(E_{+})\geq\mathbb{P}^{i}_{(S_{-},S_{+})}(E_{+})$; the next equality depends on (32) for $(S_{-},S_{+})$; and the strict inequality is due to the running cost $C_{-}$ in (3) being a sum of non-negative terms whose first, $S_{-}(1)$, is positive under $\mathbb{P}^{i}_{(S_{-},S_{+})}$. We have proved (31) and with it Proposition 4.8. ∎

### v4

@@ 3.4. Fixed-parameter ABMN asymptotics
(1,2). From $c(x)=1/\gamma(\kappa,\rho,\beta)$ and (1.8), we find that
@@
We have that $c_{-j}(x)=c\big(s_{-j}(x)\big)=1/\gamma(\kappa,\rho,\beta_{-j})$, with $\beta_{-j}$-asymptotics offered by Lemma 3.5. When $\rho=1$,
@@
$\frac{m_{-(i+1),-i}}{m_{-1,0}}$ $=$ $\prod_{j=1}^{i}\left(c_{-j}-1\right)^{-1}=\prod_{j=1}^{i}\left(\frac{1-\kappa}{1+\kappa}\right)\left(1+\frac{1-\rho}{2\rho^{2}}j^{-1}\Big(1+O(j^{-1})\Big)\right)$
@@
Noting $\big(M_{-i}/N_{-i}\big)^{\rho}=\beta_{-i}^{-\rho}\leq C/i$ offers a simplified asymptotic formula for $a_{-i}$, namely
@@
for a suitable constant $\chi=\chi(\rho,\sigma)$. From $M_{-i}=m_{-(i+1),-i}+m_{-i,-(i-1)}$ and $m_{-(i+1),-i}\ll m_{-i,-(i-1)}$, we see that
@@
Chapter 4 Nash equilibria and the ABMN equations
Here we prove Theorem 1.3 on Nash-ABMN equivalence. The forward implication $(1)\!\!\!\!\implies\!\!\!\!(2)$ is proved in the first four sections, the reverse in the fifth. The derivations follow the template given by the proof of the counterpart [24, Theorem $2.6$] in [24, Chapter $4$], with some substantial changes.
In the forward-implication proof, some arguments are new and others closely follow counterparts in [24, Chapter $4$]. To make our presentation self-contained while indicating where the overlap lies, the first three sections use the convention that Proof denotes the start of an argument with substantial new elements, while Derivation indicates one that is close to one in [24]. No lack of rigour should be inferred from use of the latter label, though we have sometimes opted for a more verbal style of presentation of such arguments. A different approach has been adopted for the reverse implication, as we explain in Section 4.5.
@@ 4.1. Escape is almost certain at a time-invariant Nash equilibrium
Recall the payoff notation (1.3). A strategy pair $(S_{-},S_{+})\in\mathcal{S}^{2}$ is said to have finite mean costs if neither $\mathbb{E}^{k}_{S_{-},S_{+}}[P_{-}]$ nor $\mathbb{E}^{k}_{S_{-},S_{+}}[P_{+}]$ equals minus infinity, for any $k\in\mathbb{Z}$.
Let $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}$. Denote by $b_{i}$ and $a_{i}$ the stakes offered by $S_{-}$ and $S_{+}$, respectively, when the counter is at site $i\in\mathbb{Z}$ (without supposing $(S_{-},S_{+})\in\mathcal{N}_{\kappa,\rho}$). The idle zone $\mathcal{I}$ is set equal to
$\mathcal{I}\,=\,\big\{j\in\mathbb{Z}:a_{j}=b_{j}=0\big\}.$
@@
Derivation. At each turn whose index lies in $\llbracket i,j-1\rrbracket$, the counter lies in the idle zone and no stakes are offered. The counter thus evolves as a symmetric simple random walk: on flip moves, by definition; on stake moves, by the zero-stake rule given in Section 1.3.1. ∎
@@
$\textrm{Rocket}_{\psi}^{\leftarrow i}(j)\,=\,\psi^{i-j+1}{\bf 1}_{j\leq i}\,\,\,,\,\,\,j\in\mathbb{Z}\,.$
@@
    $\Big(\frac{1-\kappa}{1+\kappa}\Big)^{k+1}\big(m_{\infty}-m_{*}\big)+\bigg(\psi^{k}+\Big(\frac{1-\kappa}{1+\kappa}\Big)^{k+1}\bigg)\kappa^{-1}\bigg(\frac{\psi}{1-\psi}+\frac{1-\kappa}{(1+\kappa)\big(1-\frac{1-\kappa}{(1+\kappa)\psi}\big)}\bigg)$ (4.1)
@@
    If $\textrm{Left}(S_{-})$ and $\textrm{Left}(S_{+})$ exceed -$\infty$ and $i\in\mathbb{Z}$ is less than their minimum, then, provided that the quantity given by replacing $m_{\infty}-m_{*}$ by $n_{-\infty}-n_{*}$ in (4.1) is strictly less than $n_{\infty,-\infty}$, we have that $\mathbb{E}^{i-k}_{\textrm{Rocket}_{\psi}^{\leftarrow i}(S_{-}),S_{+}}[P_{-}]>\mathbb{E}^{i-k}_{S_{-},S_{+}}[P_{-}]$.
@@
$\mathbb{E}^{i+k}_{S_{-},\textrm{Rocket}_{\psi}^{i\rightarrow}}\big[C_{+}[0,\tau_{i-1})\big]$ $=$ $\sum_{j=i}^{\infty}\mathbb{E}_{i+k}\big[\#^{[0,\tau_{i-1})}_{j}(Z)\big]\psi^{j-i+1}$ (4.2)
@@
$\mathbb{E}^{i+k}_{S_{-},\textrm{Rocket}_{\psi}^{i\rightarrow}}\big[C_{+}[\tau_{i-1},\infty)\big]\,\leq\,\big(\tfrac{1-\kappa}{1+\kappa}\big)^{k+1}\kappa^{-1}\bigg(\frac{\psi}{1-\psi}+\frac{1-\kappa}{(1+\kappa)\big(1-\frac{1-\kappa}{(1+\kappa)\psi}\big)}\bigg)\,.$ (4.3)
From (4.2) and (4.3), we find that
$\mathbb{E}^{i+k}_{S_{-},\textrm{Rocket}_{\psi}^{i\rightarrow}}\big[C_{+}\big]\,\leq\,\Big(\psi^{k}+\big(\tfrac{1-\kappa}{1+\kappa}\big)^{k+1}\Big)\kappa^{-1}\bigg(\frac{\psi}{1-\psi}+\frac{1-\kappa}{(1+\kappa)\big(1-\frac{1-\kappa}{(1+\kappa)\psi}\big)}\bigg)\,.$ (4.4)
@@
while the second is at most the right-hand side of (4.4). Hence, the hypothesis on $k$ expressed in terms of (4.1) implies that $\mathbb{E}^{i+k}_{S_{-},\textrm{Rocket}_{\psi}^{i\rightarrow}}\big[P_{+}\big]-\mathbb{E}^{i+k}_{S_{-},S_{+}}\big[P_{+}\big]$ is strictly positive, as we seek to show in proving Lemma 4.4(1).
@@
Let $(S_{-}^{\prime},S_{+})$ be a left strengthening of $(S_{-},S_{+})$. For $i\in\mathbb{Z}$, there is a coupling of gameplays $X,X^{\prime}:\mathbb{N}\to\mathbb{Z}$ under $\mathbb{P}^{i}_{S_{-},S_{+}}$ such that $X^{\prime}(j)\leq X(j)$ for $j\in\mathbb{N}$ almost surely. Couplings with the evidently needed direction for the bounds exist for each of the three other variations.
Recall that $(b_{i},a_{i})$ denotes the stake-pair offered at $i\in\mathbb{Z}$ under given $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}$.
@@
    If $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}$ satisfies $\textrm{Left}(S_{-})>$-$\infty$ and $\textrm{Left}(S_{+})=$-$\infty$, let $i\in\mathbb{Z}$ satisfy $a_{i}>0$ and $b_{j}=0$ for $j\in(-\infty,i-1\rrbracket$. Then $\mathbb{P}^{i}_{S_{-},S_{+}}(E_{-})$ equals zero.
@@ 3. $(3)$
In the remaining case, $\textrm{Left}(S_{-})>$-$\infty$ and $\textrm{Left}(S_{+})=$-$\infty$. The pair $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}\cap\mathcal{N}_{\kappa,\rho}$ is non-zero, because $S_{+}$ is; it has finite mean costs by Lemma 4.7(1). Thus $\mathbb{P}^{i}_{S_{-},S_{+}}(E^{c})=0$ by Proposition 4.3. Select $i\in\mathbb{Z}$ such that $a_{i}>0$ and $b_{j}=0$ for $j\in(-\infty,i\rrbracket$. Lemma 4.7(2) implies that $\mathbb{P}^{i}_{S_{-},S_{+}}(E_{-})=0$. Thus, $\mathbb{P}^{i}_{S_{-},S_{+}}(E_{+})=1$, so that $T_{+}$ equals $m_{\infty}$ almost surely. If Maxine plays a strategy $\hat{S}_{+}$ formed from $S_{+}$ by reducing the stake she offers at $i$ by a factor of one-half, then gameplay $X:\mathbb{N}\to\mathbb{Z}$ is equal in law under $\mathbb{P}^{i}_{S_{-},S_{+}}$ and $\mathbb{P}^{i}_{S_{-},\hat{S}_{+}}$; $T_{+}=m_{\infty}$ almost surely under both laws; but Maxine’s running cost is almost surely less under $\mathbb{P}^{i}_{S_{-},\hat{S}_{+}}$ than it is under $\mathbb{P}^{i}_{S_{-},S_{+}}$, because the first cost, incurred at site $i$, is lower. Thus, $\mathbb{E}^{i}_{S_{-},\hat{S}_{+}}[P_{+}]>\mathbb{E}^{i}_{S_{-},S_{+}}[P_{+}]$.
@@ 4.2. A Nash component wins against zero
The presence of flip moves, when $\kappa\in(0,1)$, makes the proposition non-trivial, as we now explain. In the setup in question, $S_{-}$ is known to be wide by Lemma 4.7(3); so Mina offers positive stakes at an infinite set $K$ of integer sites. When $\kappa$ equals one (as it is in [24]), so that every move is stake, this is enough to reach the desired conclusion that left escape $E_{-}$ is almost certain starting from given $i\in\mathbb{Z}$. Indeed, when $X$ visits $K$, a left move is assured; while at sites in $\mathbb{Z}\setminus K$, no stakes are offered by either player, and the next move has equal chance of being left or right, according to the rule for zero stakes given in Section 1.3.1. It is easily seen that this dynamics forces the counter leftward, through a sequence of one-way locks. However, when $\kappa\in(0,1)$, flip moves occur with probability $1-\kappa$; so, when $X$ visits $K$, the next move is left with probability $(1+\kappa)/2$. The counter thus evolves as a symmetric simple random walk on $\mathbb{Z}\setminus K$, with moves biased to the left by a uniform amount on visits to $K$. Although $K$ is infinite (since $S_{-}$ is wide), this set could in principle be arbitrarily sparse; in which case, this dynamics will not realize left escape $E_{-}$ for some (or indeed all) starting points.
@@
We will in fact prove the stronger assertion that, when $(S_{-},S_{+})\in\mathcal{N}_{\kappa,\rho}$ with $S_{-}\in\mathcal{S}_{0}$ satisfies $\mathbb{P}^{i}_{(S_{-},0)}(E_{-})=0$ for some $i\in\mathbb{Z}$ such that Mina’s first-turn stake $S_{-}(0)$ is strictly positive when $X_{0}=i$, then $(S_{-},S_{+})\not\in\mathcal{N}_{\kappa,\rho}$. Fixing such an $i$, we will show that
$\mathbb{E}^{i}_{(0,S_{+})}[P_{-}]>\mathbb{E}^{i}_{(S_{-},S_{+})}[P_{-}]\,:$ (4.5)
it is in Mina’s interests to play the zero strategy, rather than $S_{-}$, against Maxine’s $S_{+}$, when play starts at $i$. Naturally, (4.5) implies that $(S_{-},S_{+})\not\in\mathcal{N}_{\kappa,\rho}$, so proving (4.5) is enough.
Preparing to show (4.5), note that
$\mathbb{P}^{i}_{(S_{-},S_{+})}(E_{-})=\mathbb{P}^{i}_{(0,S_{+})}(E_{-})=0\,.$ (4.6)
Indeed, $(S_{-},0)\to(S_{-},S_{+})$ is a right strengthening and $(S_{-},S_{+})\to(0,S_{+})$ is a left weakening, so (4.6) follows from $\mathbb{P}^{i}_{(S_{-},0)}(E_{-})=0$ and Lemma 4.6.
Why may we expect (4.5) to hold? In other words, why would Mina switch from $S_{-}$ to 0 against $S_{+}$? That $\mathbb{P}^{i}_{(S_{-},S_{+})}(E_{-})$ is zero makes Mina’s motivation simple: $S_{-}$ is not working out for her, because her victory $E_{-}$ never happens. By switching to 0, she will save on running costs. As for terminal receipts, these are split between non-escape $E^{c}$ and right escape $E_{+}$ when she plays $S_{-}$. By playing 0 instead, Mina will cease to exert any left pressure, so, in an instance of right strengthening and monotonicity. any change to this split will take the form of a rightward move of probability mass from $E^{c}$ to $E_{+}$. But that would help Mina, because $E^{c}$ is the worse outcome for her in the sense that $n_{*}\leq n_{\infty}$. To record these inferences symbolically,
@@
where the first equality is due to absence of running cost for Mina when she plays zero; the second equality crucially invokes $\mathbb{P}^{i}_{(S_{-},0)}(E_{-})=0$; the first inequality is due to the (4.6)-consequences
@@
and the monotonicity deduction $\mathbb{P}^{i}_{(0,S_{+})}(E_{+})\geq\mathbb{P}^{i}_{(S_{-},S_{+})}(E_{+})$; the next equality depends on (4.6) for $(S_{-},S_{+})$; and the strict inequality is due to the running cost $C_{-}$ in (1.2) being a sum of non-negative terms whose first, $S_{-}(0)$, is positive under $\mathbb{P}^{i}_{(S_{-},S_{+})}$. We have proved (4.5) and with it Proposition 4.8. ∎